Find the EAR for the stated rate (APR) of 12.4 percent compounded quarterly. (Enter your answer as a percentage, omit the "%" sign in your response, and round your answer to 2 decimal places. For example, 0.12345 or 12.345% should be entered as 12.35.)
Answer #1

Effective annual rate is the rate of interest an investor earns in a year after accounting for the effects of compounding.

It is calculated using the below formula:

EAR= (1+i/n)^n-1

Where i is the interest rate and n is the number of compounding periods in one year.

EAR= (1+0.124/4)^4-1

        = 1.1299-1

        = 0.1299*100= 12.99%

Therefore, the effective annual rate is 12.99%.